What is the rate of heat conduction through the 4.11-cm-thick fur of a large animal having a 1.9-m2 surface area? Assume that the animal’s skin temperature is 33.99 ºC, that the air temperature is -4.34ºC, and that fur has the same thermal conductivity as air


1
Expert's answer
2021-03-31T09:40:29-0400
"P=\frac{kA(T_h-T_c)}{d}\\
P=\frac{(0.02435)(1.9)(33.99-(-4.34))}{0.0411}=43.15\ W"